OpenStudy (anonymous):

Which statement describes the translation of the graph of y=4(x-4)^2-2 from standard position? Moved up and to the right. Moved down and to the right. Moved down and to the left. Move up and to the left.

OpenStudy (across):

If you plot this, then you'll answer this question yourself. :)

OpenStudy (anonymous):

Okay but how will it be a translation??

OpenStudy (anonymous):

okay i plotted the parabola where would the translation be? thats what I'm confused about..

OpenStudy (across):

Translation simply means movement. In other words, what's the difference between \(y=x^2\) and \((x-4)^2-2\)? In this case, both horizontal and vertical translations are going on. How much does the graph move to the left/right? How much does it move up/down?

OpenStudy (anonymous):

okay but from the vertex? What's the starting point?? I'm so confused...

OpenStudy (across):

It doesn't matter what the starting point is because the system is the two-dimensional euclidean plane; for simplicity, you could pick the vertex.

OpenStudy (anonymous):

Well it moves and up and to the left and right..so the starting point or point you're moving from would determine the answer??

OpenStudy (across):

Think of the graph as a whole, not a single point in it. How is \(y=x^2\) different from \(y=(x-4)^2-2\)? Here, I will draw it for you: